OpenStudy (anonymous):

two bags each contain blue marbles and black marbles. the first contains 4 blue marbles and 6 black marbles and the second bag contains 6 blue marbles and 6 black marbles. a marble is randomly drawn from each bag. what is the probability that both marbles are blue? a. 1/5 b. 1/12 c. 3/10 d. 1/10

OpenStudy (ivettef365):

the first bag have 4 blues and 6 blacks, so probability would be 4/10 = 2/5 the second bag have 6 blues and 6 blacks, so probability would be 6/12 = 1/2 multiply 2/5 . 1/2

OpenStudy (anonymous):

i got 0.2 .-.

OpenStudy (ivettef365):

2 1 2 1 -- x --- = ----- = ---- 5 2 10 5

OpenStudy (ivettef365):

.2 is the same as 1/5

OpenStudy (anonymous):

thanks, :) since your already helping me with this one can you help me with another one of these? please? :c

OpenStudy (ivettef365):

sure

OpenStudy (anonymous):

two bags each contain white marbles and yellow marbles. the first contains 6 white marbles and 5 yellow marbles and the second bag contains 4 white marbles and 3 yellow marbles. a marble is randomly drawn from each bag. what is the probability that both are yellow? a. 15/77 b. 18/33 c. 10/77 d. 24/77

OpenStudy (anonymous):

i think it's A because i did exactly what you did on the other ;P

OpenStudy (ivettef365):

that is a similar questions, you have first bag = 6 white and 5 yellows = total altogether = 11 second bag: 4 white and 3 yellows: total 7 so probability of yellow on first bag 5/11 second bag = 3/7

OpenStudy (ivettef365):

now multiply both fractions

OpenStudy (anonymous):

i got it, thanks :D

OpenStudy (ivettef365):

yw :)

OpenStudy (anonymous):

34 brown marbles and a big purple bowling ball

OpenStudy (anonymous):

The probability of drawing 3 red marbles in a row from a bag of 3 blue and 4 red marbles without replacement is right at about 0.11 or about 11 in 100, if you want odds. Remember that probability is a pure number between zero (no chance - it's impossible) and 1 (it will happen - it's a certainty). A probability of 0.5 is a 50-50 shot, like a coin toss. We good? Let's move on. We have 7 marbles, and there is a 4 in 7 shot of getting a red marble in the first draw. There is a 3 in 6 shot the next draw since now we only have 6 marbles total (NO replacement) and only 3 red ones (we already drew a red one out). Similarly, there is a 2 in 5 shot of getting a third red one if the first two are red. The probability of the first action (getting a red marble on the first draw) is 4/7 or a probability of about 0.5714. Second draw, it's 3/6 or a probability of 0.5. Last, it's 2/5 or a probability of 0.40. To find the intersection of these probabilities, we just multiply them all together, and we'll get right at about 0.011 for a probability. Note: There are comments associated with